# No Language is an Island: Unifying Chinese and English in Financial Large Language Models, Instruction Data, and Benchmarks

### v1

@@ 7.1 Evaluation Setup
In addition to the FinLLMs we mentioned earlier, such as CFGPT, DISC-FinLLM, FinMA, we also select the representative and outstanding LLMs for evaluating performance. 1) ChatGPT (OpenAI   [^9] Https://www.openai.com/chatgpt): ChatGPT-3.5 and ChatGPT-4, developed by OpenAI, are widely recognized LLMs with versatile functionalities. 2) LLaMa (Touvron et al., 2023): LLaMa and advanced LlaMa2 are powerful LLMs developed by Meta AI, with parameter sizes ranging from 7 billion to 650 billion. 3) Baichuan (Yang et al., 2023a): It is a LLM developed by Baichuan-inc for Chinese and English NLP tasks, available in Baichuan-7B and Baichuan-7B-Chat variants. 4) ChatGLM  (Du et al., 2021): It is an advanced bilingual LLM developed by Tsinghua University and ZhiPu AI, with versions like ChatGLM2-6B and ChatGLM3-6B. 5) BLOOM  (Muennighoff et al., 2022): It developed by the BigScience team, is a versatile LLM with parameter sizes ranging from 560M to 176B, including Bloomz-7b1. 6) InternLM (Team, 2023): InternLM is a language model developed by SenseTime, featuring different parameter configurations such as InternLM-7B. 7) Qwen (Bai et al., 2023): It is a series of advanced LLMs introduced by Alibaba Cloud, including Qwen-7B for conversational capabilities in Chinese.
@@
Following  (Wu et al., 2023; Li et al., 2023b; Xie et al., 2023b), we employed two common used zero-shot and few-shot evaluation methods to assess the model’s adaptability and performance.
7.2 Results
We present evaluation results of different LLMs, FinLLMs and ICE-INTERN variants in Figure 4.
As shown in Figure 4, A demonstrates ICE-INTERN’s outstanding performance across various specific tasks, particularly excelling in tasks involving DLC, DLE, and DTT data types, showcasing strong Chinese language capabilities. However, in D-E, it slightly lags behind ChatGPT-4 in English and DOF tasks. It is particularly evident in F. In G, ICE-INTERN excels in 10 Chinese-English bilingual NLP tasks , except in English reasoning and DOF tasks. H further confirms ICE-INTERN ’s overall superiority in handling tasks in both English and Chinese, particularly excelling in Chinese.
@@ 7.2 Results
Model Overall Performance. Among the total of 63 specific-task metrics, ICE-INTERN variants achieves optimal performance in approximately 68%. Among them, ICE-INTERN-Full, which fine-tuned on the entire dataset, exhibits the highest percentage of all tasks with the best performance. This fully demonstrates the exceptional superiority of ICE-INTERN and highlights the importance of domain instruction fine-tuning in improving the performance of FinLLMs. InternLM overall outperforms LLaMA2, which fully validates our choice of using InternLM instead of LLaMA as the backbone model used in similar work  (Yang et al., 2023c; Wu et al., 2023; Xie et al., 2023b). We observe that that DLC has little impact for ICE-INTERN-dlc-7B. New introduction of DLE further improves the performance of ICE-INTERN-dle-7B significantly. Conversely, introducing DTT for fine-tuning ICE-INTERN-dtt-7B does not result in an improvement and instead leads to decrease. Finally, incorporating DTE for fine-tuning ICE-INTERN-Full-7B demonstrates a substantial improvement. Our discoveries align with prior research (Chen et al., 2023; Xie et al., 2023b), illustrating the efficacy of the InternLM backbone in DLC. Additionally, diverse DLE data types encompassing comprehension and reasoning tasks prove essential for optimizing performance. Moreover, our results underscore the detrimental impact of poor-quality translation datasets on fine-tuned model efficacy. These insights pave the way for innovative advancements in LLM research within the finance domain.
Bilingual Proficienc Analysis. In the cross-lingual tasks of FinHC, FinSA, and FinSP, the ICE-INTERN-Full model exhibits performance improvements ranging from approximately 1% to 10% when incorporating the corresponding translation data in fine-tuning LLMs for the same task. These findings suggest that the impact of supplementing the corresponding translation data on fine-tuning results is not as pronounced for shorter-text tasks (such as FinSA and FinHC) in DLC. However, when dealing with the long-text (including time-series data) task of FinSP in DLC, incorporating the corresponding translation data can significantly enhance the results. This is because short-text specific tasks involves relatively simple grammar and structure of the financial descriptions, and our ICE-INTERN models can already utilize sufficient information in Chinese for answer reasoning. However, when dealing with longer and more complex texts such as those containing time, historical stock prices, and tweets, our ICE-INTERN models benefit from the inclusion of English data, which enhances its ability to reason with contextual information. These observations highlight the importance of considering the length and complexity of the text when determining the effectiveness of incorporating translation data for fine-tuning LLMs in cross-lingual tasks.
Defect Case Analysis. Despite the strengths exhibited by our ICE-INTERN models on known tasks, they under perform relative to ChatGPT-4 on unseen tasks like FinDC, FINER, and FinST in DOF data types, which have not been fine-tuned. This performance gap suggests that a more diverse set of domain-specific tasks is needed for effective fine-tuning. Specifically, in FinST tasks, we utilize a sophisticated prompt design that involves generating label sequences directly. However, our fine-tuned models consistently struggle to produce outputs in the desired formats. Even models like larger-parameter ChatGPT-3.5 and ChatGPT-4, which demonstrate some capability for token labeling in FinER, face difficulties in generating labels and summaries for ECTSUM and EDTSUM in FinST, especially when confronted with longer contextual information. Our conclusion aligns with PIXIU, suggesting that for unfamiliar tasks within current FinLLMs, it will be imperative to furnish more extensive domain-specific tasks and instructional data in the future to further enhance performance.
@@ 8 Limitations
While this study has positive aspects, we acknowledge limitations: 1) Parameter Constraints: Limited resources hinder training models with over 7B parameters, impacting performance on large datasets. 2) Backbone LLM Adaptability: Fine-tuned model performance varies based on backbone model adaptability to specific tasks, requiring task-specific selection. 3) Prompt Quality Variability: Even with expert annotation, instruction prompts may perform differently across models, affecting model assessment. 4) Risk of Misuse: Open-source models support research but commercial misuse poses financial risks. Regulation is crucial for responsible usage and risk mitigation.
@@ A Examples of Context and Answer in Financial Task
Table 3-5 presents the examples of context description and annotated answer in various financial task in Chinese, English and translation datatsets.
FE
Context: 国检集团和华测检测哪个比较漂亮?业务都差不多,都是检测,并且国检持有碳交易股份,两个都买了。
@@
CFiQA-SA
Context: \$CERN-在50和200MA上方整合,这里是很好的长期进入点,止损位于10MA以下-目标区域\$70。
@@
FPB
Context: Hearst will be able to consolidate about 20% of all Russian market for advertising in press after the purchase.
@@
FiQA-SA
Context: \$SLV-4.44% at 18 now AWFUL, down from 42.50
@@
Headlines
Context: Gold holds near 3-1/2 week low as investors opt for riskier assets.
@@
CConFinQA
Context: 以下是比较2007年和2008年净收入变动的分析,金额(以百万计)。<table class=’wikitable’><tr><td>1</td><td></td><td>金额(以百万计)</td> </tr> <tr> <td>2</td ><td>2007年净收入</td> <td>\$991.1</td> </tr> <tr><td>3</td> <td>零售电价</td> <td>-17.1(17.1)</td> </tr> <tr> <td>4</td> <td>购买的电力容量</td> <td>-12.0(12.0)</td> </tr> <tr> <td>5</td> <td>净批发收入</td> <td>-7.4(7.4)</td> </tr> <tr> <td>6</td> <td>其他</td> <td>4.6</td> </tr> <tr> <td>7</td><td>2008年净收入</td> <td>\$959.2</td> </tr> </table>。
@@
EnQA
Context: The following table provides a comparison of the accumulated total return of citi common stock over a period of five years, which is listed on the nyse under the ticker symbol 201cc201d. date, citi, s&p 500, s&p financials. 31-dec-2011, 100.0, 100.0, 100.0; 31-dec-2012, 150.6, 116.0, 128.8; 31-dec-2013, 198.5, 153.6, 174.7; 31-dec-2014,206.3, 174.6, 201.3; 31-dec-2015, 197.8, 177.0, 198.2; 31-dec-2016, 229.3, 198.2, 243.4.
@@
ConFinQA
Context: The following is an analysis comparing the changes in net income between 2007 and 2008. <table class=’wikitable’><tr> <td>1</td> <td> </td> <td> amount (in millions)</td> </tr> <tr><td>2</td> <td>2007 net revenue</td> <td>\$991.1</td> </tr> <tr> <td>3</td><td>retail electric price</td> <td>-17.1 ( 17.1 )</td> </tr> <tr> <td>4</td> <td>purchased power capacity</td> <td>-12.0 ( 12.0 )</td> </tr> <tr> <td>5</td> <td>net wholesale revenue</td><td>-7.4 (7.4)</td> </tr> <tr> <td>6</td> <td>other</td> <td>4.6</td></tr> <tr> <td>7</td> <td>2008 net revenue</td> <td>\$959.2</td> </tr> </table>。
@@
CACL18
Context: \$codi近期的走向趋势如下,open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30;2015-09-16, -0.8, 0.2, -1.3, 1.0, 1.0, -0.9, -0.1, -0.1, -0.3, 0.0, 0.5; 2015-09-17,-1.4, 0.9, -1.7, 1.6, 1.6, -2.0, -1.7, -1.5, -1.8, -1.6, -1.2; 2015-09-18, -0.9, 0.2, -1.5, 0.1, 0.1, -1.6, -1.8, -1.5, -1.8, -1.7, -1.3; 2015-09-21, -0.5, 0.3, -1.8, 0.7, 0.7, -1.5, -2.3, -2.1, -2.4, -2.3, -2.1; 2015 -09-22, 1.3, 3.4, -0.4, -3.1,-3.1,1.7,0.8,0.9,0.9,0.7,1.0;2015-09-23,-0.8,0.7,-0.8,0.7,0.7,1.0, 0.1,0.3,0.2,-0.0,0.2; 2015-09-24,-0.1,1.5,-0.7,0.1,0.1,0.7,0.1,0.1,0.2,-0.1,0.1; 2015-09-25,1.5, 1.9,0.0,-0.9, -0.9,1.0, 1.0, 0.8, 1.0, 0.8, 0.9; 2015-09-28, 1.2, 1.2, -0.8, -0.2, -0.2, 0.5, 1.2, 1.0, 1.2, 0.9, 1.0; 2015-09-29, 1.3, 2.1, -0.3, -1.8, -1.8, 2.0, 3.0, 2.6, 2.8, 2.9, 2.8. \$codi评级为买入, 18.6%华尔街分析师根据平均评级/目标得出的上涨。
@@
CCIKM18
Context: \$chk近期的走向趋势如下,date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2017-10-05,-0.7,0.7,-0.9,0.7,0.7,0.0,0.6,-1.3,-3.8,-5.6, -7.0; 2017-10-06, 1.4,1.7,-1.9, -2.3, -2.3, 2.0, 2.9, 1.4, -0.8,-2.8,-4.5; 2017-10-09,0.2,0.7, -1.4,-0.2, -0.2, 1.4, 2.7, 1.9,0.1,-2.2,-3.9; 2017-10-10,7.9, 8.2, 0.0, -6.7, -6.7, 6.8, 8.9, 9.0, 7.5, 5.0,3.2; 2017-10-11, -0.8, 0.8, -4.8, 1.0, 1.0, 4.1, 6.6, 7.4, 6.4, 4.2, 2.4; 2017-10-12,1.3,2.1,-1.8,-2.5,-2.5,4.4,8.1,9.7,8.8,6.8,5.3; 2017-10-13, 0.5,1.3,-0.8, 0.8, 0.8,2.0,6.2,8.2,7.8,6.3,4.6。
@@
CBigData22
Context: \$msft近期的走向趋势如下,date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2020-10-09, -2.1, 0.0, -2.1, 2.5, 2.5, -2.5, -2.8, -3.6, -4.0, -4.0,-2.5 ; 2020-10-12, -1.2, 1.1, -2.1, 2.6, 2.6, -3.9, -4.7, -5.5, -6.1, -6.3, -5.1; 2020-10-13,-0.1,1.1,-1.1,0.7,0.7,-3.0, -4.6, -5.6, -6.4, -6.5, -5.8; 2020-10-14, 1.0, 1.5, -0.8, -0.9, -0.9, -1.2, -3.3,- 4.2, -5.2, -5.5, -5.0; 2020-10-15,-1.2, 0.3, -1.7, -0.5, -0.5,0.2,-2.4,-3.2,-4.3,-4.7,-4.7; 2020-10-16, 0.2, 1.2, -0.2, 0.0, 0.0, 0.6, -1.8, -2.8, -3.8, -4.5, -4.6; 2020-10-19, 2.9, 3.8, -0.2, -2.5, -2.5, 2.4, 0.9, -0.2, -1.1, -1.9, -2.2; 2020-10-20, 0.5, 1.3, -0.7, 0.2, 0.2, 1.5, 1.1, -0.2, -1.2, -2.0, -2.2; 2020-10-21,-0.8, 1.0, -0.8, 0.1, 0.1, 0.8, 1.2, -0.1, -0.9, -1.8, -2.2; 2020-10-22, -0.4, 0.5, -1.5, 0.0, 0.0, 0.4, 1.4, -0.1, -0.7, -1.7, -2.1。
@@
ACL18
Context: The recent trend of \$codi is as follows. date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2015-03-30, 0.0, 0.4, -0.9, 0.0, 0.0, -0.0, -0.8, -1.6, -2.1, -2.4, -2.6, 2015-03-31, 0.1, 0.3, -0.8, -0.3, -0.3, 0.2, -0.4, -1.0, -1.7, -2.0, -2.2; 2015-04-01, 1.1, 1.1, -0.7, -1.2, -1.2, 1.2, 0.9, 0.5, -0.4, -0.7, -0.9; 2015-04-02, -0.2, 0.4, -0.4, -0.2, -0.2, 1.0, 1.2, 0.7, -0.1, -0.4, -0.6; 2015-04-06, -0.6, 0.5, -0.9, 0.6, 0.6, 0.2, 0.6, 0.2, -0.6, -0.8, -1.1;2015-04-07, -0.4, 1.4, -0.4, 0.2, 0.2, -0.2, 0.4, 0.0, -0.5, -1.0, -1.2; 2015-04-08, 0.0, 0.2, -0.4, 0.1, 0.1, -0.4, 0.2, 0.0, -0.4,- 1.0, -1.2; 2015-04-09, -0.5, 0.2, -1.1, 0.6, 0.6, -0.7, -0.4, -0.5, -0.7, -1.4, -1.7; 2015-04-10, 0.1, 0.4, -0.5, 0.2, 0.2, -0.5, -0.6, -0.5,-0.8, -1.5, -1.7;2015-04-13, 1.4, 1.9, 0.0, -0.7, -0.7, 0.2, 0.1, 0.3, -0.0, -0.6, -0.9. \$codi - current report filing (8-k) 2015-04-13: william blair starts compass diversified \$codi at outperform.
@@
CIKM18
Context: The recent trend of \$aal is as follows. date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2017-01-17, 1.7, 2.1, -0.2, -1.9, -1.9, 2.5, 0.9, 1.3, 2.0, 2.3, 2.0; 2017-01-18, -0.5, 0.3, -2.1, 1.9, 1.9, 0.2, -0.7, -0.7, 0.1, 0.2, 0.2; 2017-01-19, 0.8, 1.8, -0.8, -0.8, -0.8, 0.5, 0.2, -0.1, 0.8, 1.0, 1.1; 2017-01-20, -1.0, 0.3, -1.6, 1.6, 1.6, -1.1, -0.9, -1.6, -0.9, -0.6, -0.3; 2017-01-23, 2.0, 2.4, -0.4, -2.2, -2.2,0.8,1.5,0.6,1.1,1.6,1.8 2017-01-24, -1.0, 0.5, -1.6, 1.3, 1.3, -0.1, 0.3, -0.5, -0.3, 0.2, 0.5; 2017-01-25,-0.0, 0.6, -0.6, 0.8, 0.8, -0.8, -0.6, -1.1, -1.1, -0.5, -0.4; 2017-01-26, -2.3, 0.2, -2.5, 3.5, 3.5, -3.2, -3.7, -4.1, -4.4, -3.8, -3.7 2017-01-27,6.5,6.5,-0.5,-5.3,-5.3, 1.8, 1.4, 1.5, 0.9, 1.4, 1.7; 2017-01-30, 1.6, 2.3, -2.3, -4.4, -4.4, 5.5, 5.5, 5.9, 5.3, 5.6, 6.1.

### v2

@@ 7.1 Evaluation Setup
In addition to the FinLLMs we mentioned earlier, such as CFGPT, DISC-FinLLM, FinMA, we also select the representative and outstanding LLMs for evaluating performance. 1) ChatGPT (OpenAI Brown et al. (2020)): ChatGPT-3.5-Trubo and GPT4, developed by OpenAI, are widely recognized LLMs with versatile functionalities. 2) LLaMa Touvron et al. (2023): LLaMa and advanced LlaMa2 are powerful LLMs developed by Meta AI, with parameter sizes ranging from 7 billion to 650 billion. 3) Baichuan Yang et al. (2023a): It is a LLM developed by Baichuan-inc for Chinese and English NLP tasks, available in Baichuan-7B and Baichuan-7B-Chat variants. 4) ChatGLM  Du et al. (2021): It is an advanced bilingual LLM developed by Tsinghua University and ZhiPu AI, with versions like ChatGLM2-6B and ChatGLM3-6B. 5) BLOOM  Muennighoff et al. (2022): It developed by the BigScience team, is a versatile LLM with parameter sizes ranging from 560M to 176B, including Bloomz-7B1. 6) InternLM Team (2023): InternLM is a language model developed by SenseTime, featuring different parameter configurations such as InternLM-7B. 7) Qwen Bai et al. (2023): It is a series of advanced LLMs introduced by Alibaba Cloud, including Qwen-7B for conversational capabilities in Chinese.
Following  Li et al. (2023b); Xie et al. (2023b), we use two common used zero-shot and few-shot methods to evaluate model performance.
7.2 Results
(1) Benchmark Evaluation Performance The bilingual ICE-FLARE benchmark results are shown in Table 3. As shown in Table 3, the performance of ICE-INTERN and its variants in monolingual settings significantly favors Chinese over English, both across various tasks and in overall average performance. In English, the overall average performance of fully fine-tuned ICE-INTERN-full is slightly weaker than that of ChatGPT. However, across individual tasks, except for ChatGPT, ICE-INTERN and its variants perform best in the CLS and PRE NLP tasks, while their performance in the REA and OFT NLP tasks is not particularly remarkable, sometimes even weaker than other models such as FinMA and Bloomz-7B1. This indicates that full fine-tuning ICE-INTERN-full leads to a substantial overall improvement, offsetting some subpar task performances. Furthermore, we observed that apart from GPT4, fully fine-tuned ICE-INTERN-full demonstrates superior performance in bilingual tasks, particularly evident in the PRE and EXT tasks. (2) Model Balanced Performance As shown in Figure 4, A demonstrates that ICE-INTERN excels in specific tasks, particularly those involving DLC, DLE, and DTT data types, showcasing strong Chinese language capabilities. However, in D-E, it slightly lags behind GPT4 in English and DOF tasks, a distinction notably evident in F. Within G, ICE-INTERN excels in 10 Chinese-English bilingual NLP tasks, except in English reasoning and DOF tasks. H further confirms ICE-INTERN ’s overall superiority in handling tasks in both English and Chinese, particularly excelling in Chinese.
@@ 7.2 Results
(3) Model Detail Performance Among the total of 63 specific-task metrics, ICE-INTERN variants achieves optimal performance in approximately 68%. Among them, ICE-INTERN-Full, which fine-tuned on the entire dataset, exhibits the highest percentage of all tasks with the best performance. This fully demonstrates the exceptional superiority of ICE-INTERN and highlights the importance of domain instruction fine-tuning in improving the performance of FinLLMs. InternLM overall outperforms LLaMA2, which fully validates our choice of using InternLM instead of LLaMA2 as the backbone model used in similar work Yang et al. (2023c); Wu et al. (2023); Xie et al. (2023b). We observe that that DLC has little impact for ICE-INTERN-dlc-7B. New introduction of DLE further improves the performance of ICE-INTERN-dle-7B significantly. Conversely, introducing DTT for fine-tuning ICE-INTERN-dtt-7B does not result in an improvement and instead leads to decrease. Finally, incorporating DTE for fine-tuning ICE-INTERN-Full-7B demonstrates a substantial improvement. Our discoveries align with prior research Chen et al. (2023); Xie et al. (2023b), illustrating the efficacy of the InternLM backbone in DLC. Additionally, diverse DLE data types encompassing comprehension and reasoning tasks prove essential for optimizing performance. Moreover, our results underscore the detrimental impact of poor-quality translation datasets on fine-tuned model efficacy. These insights pave the way for innovative advancements in LLM research within the finance domain. (4) Bilingual Proficienc Analysis In the cross-lingual tasks of FinHC, FinSA, and FinSP, the ICE-INTERN-Full model exhibits performance improvements ranging from approximately 1% to 10% when incorporating the corresponding translation data in fine-tuning LLMs for the same task. These findings suggest that the impact of supplementing the corresponding translation data on fine-tuning results is not as pronounced for shorter-text tasks (such as FinSA and FinHC) in DLC. However, when dealing with the long-text (including time-series data) task of FinSP in DLC, incorporating the corresponding translation data can significantly enhance the results. This is because short-text specific tasks involves relatively simple grammar and structure of the financial descriptions, and our ICE-INTERN models can already utilize sufficient information in Chinese for answer reasoning. However, when dealing with longer and more complex texts such as those containing time, historical stock prices, and tweets, our ICE-INTERN models benefit from the inclusion of English data, which enhances its ability to reason with contextual information. These observations highlight the importance of considering the length and complexity of the text when determining the effectiveness of incorporating translation data for fine-tuning LLMs in cross-lingual tasks. (5) Defect Case Analysis Despite the strengths exhibited by our ICE-INTERN models on known tasks, they under perform relative to GPT4 on unseen tasks like FinDC, FINER, and FinST in DOF data types, which have not been fine-tuned. This performance gap suggests that a more diverse set of domain-specific tasks is needed for effective fine-tuning. Specifically, in FinST tasks, we utilize a sophisticated prompt design that involves generating label sequences directly. However, our fine-tuned models consistently struggle to produce outputs in the desired formats. Even models like larger-parameter ChatGPT-3.5 and GPT4, which demonstrate some capability for token labeling in FinER, face difficulties in generating labels and summaries for ECTSUM and EDTSUM in FinST, especially when confronted with longer contextual information. Our conclusion aligns with PIXIU, suggesting that for unfamiliar tasks within current FinLLMs, it will be imperative to furnish more extensive domain-specific tasks and instructional data in the future to further enhance model performance.
@@ 8 Limitations
While this study has positive aspects, we acknowledge limitations: 1) Parameter Constraints: Limited resources hinder training models with over 7B parameters, impacting performance on large datasets. 2) Backbone LLM Adaptability: Fine-tuned model performance varies based on backbone model adaptability to specific tasks, requiring task-specific selection. 3) Prompt Quality Variability: Even with expert annotation, instruction prompts may perform differently across models, affecting model assessment. 4) Risk of Misuse: Open-source models facilitate research, but commercial misuse can lead to financial risks, necessitating effective regulation.
@@ A Examples of Context and Answer in Financial Task
Table 4-6 presents the examples of context description and annotated answer in various financial task in Chinese, English and translation datatsets.
FE Context: 国检集团和华测检测哪个比较漂亮?业务都差不多,都是检测,并且国检持有碳交易股份,两个都买了。
@@
CFiQA-SA Context: \$CERN-在50和200MA上方整合,这里是很好的长期进入点,止损位于10MA以下-目标区域\$70。
@@
FPB Context: Hearst will be able to consolidate about 20% of all Russian market for advertising in press after the purchase.
@@
FiQA-SA Context: \$SLV-4.44% at 18 now AWFUL, down from 42.50
@@
Headlines Context: Gold holds near 3-1/2 week low as investors opt for riskier assets.
@@
CConFinQA Context: 以下是比较2007年和2008年净收入变动的分析,金额(以百万计)。<table class=’wikitable’><tr><td>1</td><td></td><td>金额(以百万计)</td> </tr> <tr> <td>2</td ><td>2007年净收入</td> <td>\$991.1</td> </tr> <tr><td>3</td> <td>零售电价</td> <td>-17.1(17.1)</td> </tr> <tr> <td>4</td> <td>购买的电力容量</td> <td>-12.0(12.0)</td> </tr> <tr> <td>5</td> <td>净批发收入</td> <td>-7.4(7.4)</td> </tr> <tr> <td>6</td> <td>其他</td> <td>4.6</td> </tr> <tr> <td>7</td><td>2008年净收入</td> <td>\$959.2</td> </tr> </table>。
@@
EnQA Context: The following table provides a comparison of the accumulated total return of citi common stock over a period of five years, which is listed on the nyse under the ticker symbol 201cc201d. date, citi, s&p 500, s&p financials. 31-dec-2011, 100.0, 100.0, 100.0; 31-dec-2012, 150.6, 116.0, 128.8; 31-dec-2013, 198.5, 153.6, 174.7; 31-dec-2014,206.3, 174.6, 201.3; 31-dec-2015, 197.8, 177.0, 198.2; 31-dec-2016, 229.3, 198.2, 243.4.
@@
ConFinQA Context: The following is an analysis comparing the changes in net income between 2007 and 2008. <table class=’wikitable’><tr> <td>1</td> <td> </td> <td> amount (in millions)</td> </tr> <tr><td>2</td> <td>2007 net revenue</td> <td>\$991.1</td> </tr> <tr> <td>3</td><td>retail electric price</td> <td>-17.1 ( 17.1 )</td> </tr> <tr> <td>4</td> <td>purchased power capacity</td> <td>-12.0 ( 12.0 )</td> </tr> <tr> <td>5</td> <td>net wholesale revenue</td><td>-7.4 (7.4)</td> </tr> <tr> <td>6</td> <td>other</td> <td>4.6</td></tr> <tr> <td>7</td> <td>2008 net revenue</td> <td>\$959.2</td> </tr> </table>。
@@
CACL18 Context: \$codi近期的走向趋势如下,open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30;2015-09-16, -0.8, 0.2, -1.3, 1.0, 1.0, -0.9, -0.1, -0.1, -0.3, 0.0, 0.5; 2015-09-17,-1.4, 0.9, -1.7, 1.6, 1.6, -2.0, -1.7, -1.5, -1.8, -1.6, -1.2; 2015-09-18, -0.9, 0.2, -1.5, 0.1, 0.1, -1.6, -1.8, -1.5, -1.8, -1.7, -1.3; 2015-09-21, -0.5, 0.3, -1.8, 0.7, 0.7, -1.5, -2.3, -2.1, -2.4, -2.3, -2.1; 2015 -09-22, 1.3, 3.4, -0.4, -3.1,-3.1,1.7,0.8,0.9,0.9,0.7,1.0;2015-09-23,-0.8,0.7,-0.8,0.7,0.7,1.0, 0.1,0.3,0.2,-0.0,0.2; 2015-09-24,-0.1,1.5,-0.7,0.1,0.1,0.7,0.1,0.1,0.2,-0.1,0.1; 2015-09-25,1.5, 1.9,0.0,-0.9, -0.9,1.0, 1.0, 0.8, 1.0, 0.8, 0.9; 2015-09-28, 1.2, 1.2, -0.8, -0.2, -0.2, 0.5, 1.2, 1.0, 1.2, 0.9, 1.0; 2015-09-29, 1.3, 2.1, -0.3, -1.8, -1.8, 2.0, 3.0, 2.6, 2.8, 2.9, 2.8. \$codi评级为买入, 18.6%华尔街分析师根据平均评级/目标得出的上涨。
@@
CCIKM18 Context: \$chk近期的走向趋势如下,date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2017-10-05,-0.7,0.7,-0.9,0.7,0.7,0.0,0.6,-1.3,-3.8,-5.6, -7.0; 2017-10-06, 1.4,1.7,-1.9, -2.3, -2.3, 2.0, 2.9, 1.4, -0.8,-2.8,-4.5; 2017-10-09,0.2,0.7, -1.4,-0.2, -0.2, 1.4, 2.7, 1.9,0.1,-2.2,-3.9; 2017-10-10,7.9, 8.2, 0.0, -6.7, -6.7, 6.8, 8.9, 9.0, 7.5, 5.0,3.2; 2017-10-11, -0.8, 0.8, -4.8, 1.0, 1.0, 4.1, 6.6, 7.4, 6.4, 4.2, 2.4; 2017-10-12,1.3,2.1,-1.8,-2.5,-2.5,4.4,8.1,9.7,8.8,6.8,5.3; 2017-10-13, 0.5,1.3,-0.8, 0.8, 0.8,2.0,6.2,8.2,7.8,6.3,4.6。
@@
CBigData22 Context: \$msft近期的走向趋势如下,date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2020-10-09, -2.1, 0.0, -2.1, 2.5, 2.5, -2.5, -2.8, -3.6, -4.0, -4.0,-2.5 ; 2020-10-12, -1.2, 1.1, -2.1, 2.6, 2.6, -3.9, -4.7, -5.5, -6.1, -6.3, -5.1; 2020-10-13,-0.1,1.1,-1.1,0.7,0.7,-3.0, -4.6, -5.6, -6.4, -6.5, -5.8; 2020-10-14, 1.0, 1.5, -0.8, -0.9, -0.9, -1.2, -3.3,- 4.2, -5.2, -5.5, -5.0; 2020-10-15,-1.2, 0.3, -1.7, -0.5, -0.5,0.2,-2.4,-3.2,-4.3,-4.7,-4.7; 2020-10-16, 0.2, 1.2, -0.2, 0.0, 0.0, 0.6, -1.8, -2.8, -3.8, -4.5, -4.6; 2020-10-19, 2.9, 3.8, -0.2, -2.5, -2.5, 2.4, 0.9, -0.2, -1.1, -1.9, -2.2; 2020-10-20, 0.5, 1.3, -0.7, 0.2, 0.2, 1.5, 1.1, -0.2, -1.2, -2.0, -2.2; 2020-10-21,-0.8, 1.0, -0.8, 0.1, 0.1, 0.8, 1.2, -0.1, -0.9, -1.8, -2.2; 2020-10-22, -0.4, 0.5, -1.5, 0.0, 0.0, 0.4, 1.4, -0.1, -0.7, -1.7, -2.1。
@@
ACL18 Context: The recent trend of \$codi is as follows. date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2015-03-30, 0.0, 0.4, -0.9, 0.0, 0.0, -0.0, -0.8, -1.6, -2.1, -2.4, -2.6, 2015-03-31, 0.1, 0.3, -0.8, -0.3, -0.3, 0.2, -0.4, -1.0, -1.7, -2.0, -2.2; 2015-04-01, 1.1, 1.1, -0.7, -1.2, -1.2, 1.2, 0.9, 0.5, -0.4, -0.7, -0.9; 2015-04-02, -0.2, 0.4, -0.4, -0.2, -0.2, 1.0, 1.2, 0.7, -0.1, -0.4, -0.6; 2015-04-06, -0.6, 0.5, -0.9, 0.6, 0.6, 0.2, 0.6, 0.2, -0.6, -0.8, -1.1;2015-04-07, -0.4, 1.4, -0.4, 0.2, 0.2, -0.2, 0.4, 0.0, -0.5, -1.0, -1.2; 2015-04-08, 0.0, 0.2, -0.4, 0.1, 0.1, -0.4, 0.2, 0.0, -0.4,- 1.0, -1.2; 2015-04-09, -0.5, 0.2, -1.1, 0.6, 0.6, -0.7, -0.4, -0.5, -0.7, -1.4, -1.7; 2015-04-10, 0.1, 0.4, -0.5, 0.2, 0.2, -0.5, -0.6, -0.5,-0.8, -1.5, -1.7;2015-04-13, 1.4, 1.9, 0.0, -0.7, -0.7, 0.2, 0.1, 0.3, -0.0, -0.6, -0.9. \$codi - current report filing (8-k) 2015-04-13: william blair starts compass diversified \$codi at outperform.
@@
CIKM18 Context: The recent trend of \$aal is as follows. date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2017-01-17, 1.7, 2.1, -0.2, -1.9, -1.9, 2.5, 0.9, 1.3, 2.0, 2.3, 2.0; 2017-01-18, -0.5, 0.3, -2.1, 1.9, 1.9, 0.2, -0.7, -0.7, 0.1, 0.2, 0.2; 2017-01-19, 0.8, 1.8, -0.8, -0.8, -0.8, 0.5, 0.2, -0.1, 0.8, 1.0, 1.1; 2017-01-20, -1.0, 0.3, -1.6, 1.6, 1.6, -1.1, -0.9, -1.6, -0.9, -0.6, -0.3; 2017-01-23, 2.0, 2.4, -0.4, -2.2, -2.2,0.8,1.5,0.6,1.1,1.6,1.8 2017-01-24, -1.0, 0.5, -1.6, 1.3, 1.3, -0.1, 0.3, -0.5, -0.3, 0.2, 0.5; 2017-01-25,-0.0, 0.6, -0.6, 0.8, 0.8, -0.8, -0.6, -1.1, -1.1, -0.5, -0.4; 2017-01-26, -2.3, 0.2, -2.5, 3.5, 3.5, -3.2, -3.7, -4.1, -4.4, -3.8, -3.7 2017-01-27,6.5,6.5,-0.5,-5.3,-5.3, 1.8, 1.4, 1.5, 0.9, 1.4, 1.7; 2017-01-30, 1.6, 2.3, -2.3, -4.4, -4.4, 5.5, 5.5, 5.9, 5.3, 5.6, 6.1.